28. Slope 1/2; y intercept (o, -6)
Can you please help me with this one?
Thanks
Carolyn

Answer by tutorcecilia(2152)

The slope-intercept equation form of a line is: y=mx+b where:
y= any y-value on the line
m=the slope
x= the x-value that corresponds to the y-value of the line
b=the y-intercept
.
So,since the slope = 1/2 and the y-intercept = (0, -6) or just (-6), than
plugging these values into the formula:
y=1/2x-6 becomes the equation of the line.